# Measure Twice, Match Once: Flexible Graph Similarity Computation With A Proactive Optimization Strategy

### v1

@@ Abstract.
Graph Edit Distance (GED) is an important similarity measure in graph retrieval, which quantifies the minimum cost of transforming one graph into another through edit operations, and offers flexibility by allowing customizable operation costs. Recent learning-based approaches approximate GEDs with the distances between representations in vector spaces. However, these methods often struggle with varying operation costs due to neglecting the impact of these costs on determining optimal graph mappings. Furthermore, they rely on isolated node distances as guidance, necessitating inefficient reactive refinements of mappings. To address these issues, we propose Graph Edit Network (Gen), a novel learning-based approach for flexible GED computation. By identifying the limitations of existing methods in capturing flexibility of GED, we introduce a principled yet simple solution that incorporates the operation costs before establishing mappings. To improve matching efficiency, we propose a strategy that proactively optimizes guidance from a graph perspective. This strategy initializes guidance as each node’s alignment difficulty and captures the interdependencies between matches within and across graphs through a difficulty propagation mechanism, enabling more informed decisions. As a result, Gen selects optimal matches in a single step, minimizing the need for costly refinements. Results on real-world and synthetic datasets demonstrate the effectiveness, time efficiency, and adaptability of Gen, achieving up to 37.8% error reduction and 72.7% inference time reduction compared with state-of-the-art models, while performing robustly under varying cost settings and graph sizes.
graph similarity computation, graph edit distance, neural networks
ccs: Information systems Similarity measuresccs: Computing methodologies Neural networks
Figure 1. The black dashed lines are matches, and the red markings are edits. $f_{1}$ and $f_{2}$ are mappings. $f_{1}$ incurs 2 node substitutions (ns), 1 node insertion, 1 edge deletion, and 2 edge insertions (ei), denoted as $[2,1,0,1,2]$, while $f_{2}$ incurs $[4,1,0,0,1]$. When ns costs 2, other operations cost 1, $f_{1}$ is optimal with a cost of 8, while $f_{2}$ costs 10. However, if ei costs 2, others cost 1, $f_{1}$ costs 8 while $f_{2}$ costs 7, making $f_{2}$ the optimal choice.
1. Introduction
@@
Graph similarity computation is crucial for graph retrieval systems (Zheng et al., 2014; He and Singh, 2006; Wang et al., 2010; Zhu et al., 2012; Zheng et al., 2013; Chang et al., 2023), as it enables the discovery of relevant patterns and relationships. One key measure of graph similarity is Graph Edit Distance (GED), which quantifies the minimum cost of transforming one graph into another through operations, such as node/edge insertion, deletion, or substitution. The flexibility of GED lies in its ability to assign customizable costs to each operation, enabling similarity measures tailored to specific applications. For example, in graph database for recommender systems, higher costs can be assigned to nodes/edges representing key user preferences. This prioritizes the retrieval of graphs that preserve these preferences, which can be used to enhance personalized recommendations.
Conventionally, GED is computed using A*-like algorithms (Abu-Aisheh et al., 2015; Zeng et al., 2009; Chang et al., 2023), which sequentially establish node/edge mappings between graphs. The operation costs assigned to individual operations are crucial in this process, as they directly influence matching decisions. This leads to varying optimal mappings, ultimately affecting the final GED, as illustrated in Fig. 1. However, such a process requires exploring all possible mappings to find the optimal ones with the minimum total cost, resulting in significant latency. Recent studies show that state-of-the-art algorithms struggle to compute exact GED for small graphs with 16 nodes within a reasonable time (Blumenthal and Gamper, 2020). This limits their practicality for modern graph retrieval tasks that require the timely processing of large volumes of graphs.
Recently, with the great success of Graph Neural Networks (GNNs) (Kipf and Welling, 2017; Xu et al., 2019; Veličković et al., 2018; Brody et al., 2022), numerous GNN-based approaches have been proposed (Li et al., 2019; Qin et al., 2021; Zhuo and Tan, 2022; Ranjan et al., 2022; Bai et al., 2019, 2020; Zhang et al., 2021; Doan et al., 2021; Bai and Zhao, 2021; Liu et al., 2022; Piao et al., 2023; Jain et al., 2024; Jia et al., 2023). By leveraging GNNs to generate expressive node representations that capture rich neighborhood information, they effectively approximate GED in the embedding space. These methods are promising alternatives for graph retrieval tasks, where response speed is prioritized over exact mappings. Depending on how node representations are utilized, they can be divided into two categories: graph comparison methods and node comparison methods. Graph comparison methods (Li et al., 2019; Qin et al., 2021; Zhuo and Tan, 2022; Ranjan et al., 2022; Bai and Zhao, 2021) aggregate node representations into a single graph embedding and compute GED by measuring the distance between these embeddings. While computationally efficient, this aggregation leads to a loss of node-level details, which are crucial for considering edit operations. Consequently, these approaches may fail to capture the flexibility inherent in GED computation. In contrast, node comparison methods (Bai et al., 2019; Li et al., 2019; Bai et al., 2020; Zhang et al., 2021; Doan et al., 2021; Piao et al., 2023; Jain et al., 2024; Liu et al., 2022) compare node representations between graphs to compute node matching costs. These costs guide the iterative refinements of mappings by down-weighting high-cost matches for each node. The total cost of the final mappings is then used to approximate the GED. By focusing on node-level differences, similar to conventional algorithms, methods in this category may offer better flexibility.
However, these node comparison methods face two major limitations: (1) Neglect the Role of Operation Costs in Mapping Determination. Most methods fail to account for operation costs. GraphEdX (Jain et al., 2024) attempts to address this by initially establishing mappings without considering operation costs, and then using these costs to reweigh the total costs of mappings afterward. However, operation costs directly influence matching decisions, meaning mappings should vary according to these costs, not just the final GED value. This oversight leads to static mappings under varying cost settings, resulting in inaccurate GED predictions. (2) Reliance on Insufficient Node-level Guidance. These methods use node matching costs as guidance to establish mappings. However, these costs struggle to effectively guide the matching process because they capture only isolated node-level similarities. GED computation, in contrast, requires comparing the interdependencies of matched nodes within graphs to quantify and minimize structural mismatches between graphs after alignment. Node-level guidance lacks this necessary graph perspective. As a result, existing methods rely on a reactive mapping refinement strategy, which improves mappings over multiple rounds with additional constraints, such as considering local node connections (Li et al., 2019; Bai et al., 2020) or minimizing graph transport costs (Doan et al., 2021; Piao et al., 2023; Jain et al., 2024). These refinements significantly increase computational overhead, leading to an inefficient matching process.
To address these limitations, we propose Graph Edit Network (Gen), a novel learning-based model for flexible GED approximation. Gen introduces two key features: (1) Flexible Adaptation to Cost Settings. We identify a key limitation of existing methods in capturing the flexibility of GED: they either neglect or treat operation costs as irrelevant to mapping determination. To address this, Gen incorporates operation costs prior to establishing mappings. This solution aligns with the principle in GED computation, where mappings are dynamically determined based on operation cost, enabling flexible mapping determination across varying cost settings. (2) Effective Graph Perspective Guidance. Unlike existing methods that rely on insufficient node-level guidance and require reactive multi-round refinements to improve mappings, we propose a strategy that proactively optimizes guidance from a graph perspective, thereby minimizing the need for such refinements. First, this strategy initializes guidance as each node’s alignment difficulty based on the global cost ratios of its associated matches. Such guidance provides graph-level comparability and mitigating the risk of myopic, locally optimal decisions. Second, it uses a Difficulty Propagation mechanism to propagate these difficulties within and between graphs, capturing the interdependencies of nodes involved in matches. Within graphs, it respects the node connections, and estimates the alignment difficulty for each neighborhood, broadens the decision-making scope from individual nodes to their neighborhoods. Between graphs, it considers the neighborhood alignment difficulties of involved nodes for each match, facilitating more balanced match selections. Leveraging this guidance, Gen selects optimal matches in a single step, ensuring an efficient matching process. The main contributions are summarized as follows:
  • We propose Graph Edit Network (Gen), which enables flexible GED computation under varying cost settings and employs an effective proactive guidance optimization strategy.
  • We introduce a principled yet simple solution to model the inherent flexibility of GED computation, addressing a gap in graph similarity learning.
  • We propose a proactive optimization strategy with a difficulty propagation mechanism that provides effective graph perspective guidance, allowing optimal match selections in a single step and ensuring an efficient matching process.
Empirical results on three commonly used GED datasets validate the efficiency of the proposed method, achieving up to 37.8% error reduction and 72.7% inference time reduction compared with state-of-the-art models. Gen demonstrates robustness across varying cost settings, consistently delivering strong performance. The ablation study highlights the importance of our proposed strategy, and the hyperparameter sensitivity analysis provides insights into Gen’s behavior. Additionally, experiments on larger synthetic graphs highlight Gen’s reliable performance across different node sizes, confirming its practical utility in real-world applications.
2. Related Work
Graph similarity computation (GSC) research approximates domain-agnostic measures, such as graph edit distance (GED) (Bai et al., 2019; Li et al., 2019) or maximum common subgraph (MCS) (Roy et al., 2022), using neural networks. These GSC methods can be categorized into end-to-end approaches that rely on GNNs and hybrid models that combine GNNs with A* algorithms (Wang et al., 2021; Yang and Zou, 2021; Liu et al., 2023). This paper focuses on end-to-end GED approximation, where GSC methods fall into two main categories: graph comparison methods and node comparison methods.
Graph comparison methods rely on graph-level similarity, with representative methods including (Li et al., 2019; Qin et al., 2021; Zhuo and Tan, 2022; Ranjan et al., 2022; Bai and Zhao, 2021). GMN-emb (Li et al., 2019) generates graph embeddings using a gated mechanism to capture node importance. To improve efficiency, EGSC (Qin et al., 2021) employs a teacher network to fuse graph pair embeddings for contextual predictions, while using a student network to distill them into separate embeddings for storage. Similarly, Eric (Zhuo and Tan, 2022) incorporates a node-graph alignment constraint during training and decouples this module during inference to reduce response time. To model the triangle inequality of the metric, Greed (Ranjan et al., 2022) uses the lower and upper bounds of GED computed by inexact algorithms to frame the distances between graph representations. To improve the interpretability, TaGSim (Bai and Zhao, 2021) predicts graph edit vectors (GEVs) to represent the counts of each operation type by comparing graph representations at multiple scales. Among these methods, while TaGSim might appear capable of capturing flexible GED computation, such a graph comparison method is insensitive to mapping variations. Reweighing its GEVs based on operation costs, as in GraphEdX (Jain et al., 2024), can only lead to inaccurate GED predictions.
@@ 2. Related Work
Node comparison methods focus on node-level similarities, with representative works including (Li et al., 2019; Bai et al., 2019, 2020; Doan et al., 2021; Piao et al., 2023; Jain et al., 2024). To enhance graph similarity scores, SimGNN (Bai et al., 2019) introduces histogram features based on node similarities. However, these non-differentiable features limit the method’s learning capability, as they cannot participate in training. Except for SimGNN, other approaches adopt a reactive mapping refinement strategy. GMN-match (Li et al., 2019) repeats a fusion-propagation process, starting by fusing representations of cross-graph node pairs using matching-cost-based local attention. The fused representations are then propagated within graphs and adjusted according to local node connections. However, node-level fusion ratios can lead to myopic, locally optimal decisions that directly affect subsequent updates, increasing the risk of error propagation. GraphSim (Bai et al., 2020) reorders multi-scale similarity matrices using BFS traversal to consider local node connections. However, this traversal oversimplifies the complex graph structures, potentially leading to inconsistent performance. Recent methods adopt combinatorial optimization techniques, using the transport cost of established node mappings to approximate GED. GOTSim (Doan et al., 2021) augments the similarity matrix to account for node deletion and insertion operations, then uses a linear assignment solver to align nodes at multiple scales. GEDGNN (Piao et al., 2023) and GraphEdX (Jain et al., 2024) use Gumbel-Sinkhorn networks to iteratively refine mappings. GEDGNN computes both a similarity matrix and a matching matrix to enhance performance. GraphEdX (Jain et al., 2024) designs distance measures for each operation type to establish mappings and subsequently reweighs transport costs with operation costs. However, these optimization-based methods treat graph combinatorial optimization in a matrix formulation, overlooking the node connections. As a result, the cost of their established mappings may fail to reflect the true structural differences between graphs. Furthermore, all methods in this category overlook the role of operation costs in determining mappings, leading to static mappings under varying cost settings.
3. Background
Figure 2. Overview of our proposed Graph Edit Networks (Gen). The dashed node represents the dummy node.
@@ 3.1. Problem Setup
Given a source graph $S$ and a target graph $T$, where the graph with fewer nodes is padded with dummy nodes to ensure that $|\mathcal{V}_{S}|=|\mathcal{V}_{T}|$. The GED computation can be viewed as an error-correcting graph matching problem (Bunke, 1997), which constructs a bijective node mapping $f:\mathcal{V}_{S}\mapsto\mathcal{V}_{T}$ that minimizes the total matching cost. The solution space of such mappings can be represented as a complete bipartite graph $\mathcal{B}=(\mathcal{V}_{S}\cup\mathcal{V}_{T},\mathcal{E}_{\mathcal{B}})$, where each edge $m\in\mathcal{E}_{\mathcal{B}}$ signifies a candidate source-target node match.
Matching Cost. Elements across the graphs are matched by applying edit operations to correct the inherent differences, such as label mismatches. Such corrections incur matching costs $\text{c}(i_{s},i_{t})$, representing the operation cost-weighted differences, as follows.
(1) $\text{c}(i_{s},i_{t})=\text{cost}(op_{(i_{s},i_{t})})\cdot d(i_{s},i_{t})$
Where $\text{cost}(\cdot)$ is a non-negative, customizable cost setting function that assigns an operation cost to each edit $op_{(i_{s},i_{t})}$. $d(i_{s},i_{t})$ quantifies the attribute and structural differences between elements $i_{s}\in S$ and $i_{t}\in T$, which is typically 0 or 1 for nodes/edges, as they are individual elements. For simplicity, $d(\cdot,\cdot)$ is omitted in such cases.
Graph Edit Distance (GED). Given any node mapping $f$, the corresponding operation sequence can be easily deduced, and vice versa. The total cost of $f$ and GED can be computed as follows.
(2) $c(S,T)_{f}=\sum_{s\in\mathcal{V}_{S},t\in\mathcal{V}_{T}}\text{c}(s,t)+\sum_{e\in f(\mathcal{A}_{S})\setminus\mathcal{A}_{T}}\text{c}(e,\epsilon)+\sum_{e^{\prime}\in\mathcal{A}_{T}\setminus f(\mathcal{A}_{S})}\text{c}(\epsilon,e^{\prime})$
(3) $\text{GED}(S,T)=min(c(S,T)_{f}|f\subset\mathcal{E}_{\mathcal{B}})$
Where $\epsilon$ represents nonexistent edges, $\text{cost}(s,t)$ signifies the matching cost between $s$ and $t$, $e\in f(\mathcal{A}_{S})\setminus\mathcal{A}_{T}$ refers to an edge $e(s,s^{\prime})$ in $S$ without a correspondence in $T$, and similarly for $e^{\prime}\in\mathcal{A}_{T}\setminus f(\mathcal{A}_{S})$.
Limitation for Flexible GED Computation. When $\text{cost}(\cdot)$ in Eq. 1 assigns a uniform cost (e.g., 1), to each operation, the total cost corresponds to the number of applied operations. Suppose $f^{*}$ is an optimal mapping under this setting. When $\text{cost}(\cdot)$ varies, $c(S,T)_{f^{*}}$ changes accordingly, causing previously optimal mappings to become suboptimal, as illustrated in Fig. 1. Existing methods either neglect or treat operation costs as irrelevant when determining the mappings, resulting in static mappings based solely on inherent element differences and failing to capture the flexibility of GED. Consequently, these methods are incapable of performing customizable searches that adapt to varying cost settings. Furthermore, their reliance on static mappings constrains both theoretical and practical understanding of GED computation.
Problem Formulation. We use the total matching cost $c(S,T)$ of the soft node mapping to approximate GED values, which ranges from $[0,+\infty)$. The total mapping cost is distributed across selected matches. Specifically, each candidate match $m$ in $\mathcal{B}$ is associated with a matching cost, reflecting the cumulative cost needed to transform $\mathcal{N}(s)$ into $\mathcal{N}(t)$. A compensation approximator $g(\cdot)$ is introduced to adjust for duplicated edge cost calculations caused by overlapping neighborhoods. Furthermore, each match is assigned a certainty $\mathbf{1}_{s,t}$ to indicate its soft selection or rejection. The total mapping cost is then expressed as follows.
(4) $c(s,t)=g\bigg{(}\sum_{op\in|\text{op}|}cost(op)\cdot d(\mathcal{N}(s),\mathcal{N}(t))_{op}\bigg{)}$
(5) $c(S,T)=\sum_{s\in|\mathcal{V}_{S}|,t\in|\mathcal{V}_{T}|}\mathbf{1}_{s,t}\cdot c(s,t)$
Where $c(s,t)$ is the matching cost, $cost(op)$ is the operation cost for $op$ applied between $s$ and $t$. For each type $\in\text{op}$ op, $d(\cdot,\cdot)_{op}$ quantifies the differences between $\mathcal{N}(s)$ and $\mathcal{N}(t)$ corrected by $op$.
4. The Proposed GEN Architecture
The overview of our proposed Gen is illustrated in Figure 2. Given a source-target graph pair and customized cost per operation type for each possible match as input, Gen begins by generating node representations for both graphs using a shared GNN Encoder. To enable flexible GED computation, Gen dynamically incorporates operation costs with node distances within the embedding space during the Node Matching Cost Estimation step. Next, Gen proactively optimizes guidance from a graph perspective to minimize the need for reactive refinements. It begins by initializing the guidance as each node’s alignment difficulty to enable globally comparability. Specifically, it first normalizes the matching costs across all potential node matches as global cost ratios, then computes each node’s alignment difficulty as the total cost ratios of its associated matches. To capture the interdependencies of matches, it propagates these difficulties within and between graphs through the Difficulty Propagation mechanism. Within each graph, it respects node connections and estimates the alignment difficulty of each node’s neighborhood. Between graphs, Gen calculates the certainty for each match by considering the neighborhood alignment difficulties of its involved nodes, thus balancing match selection. Finally, the overall matching cost of a graph is determined by summing the certainty-weighted matching costs.
4.1. GNN Encoder
(6) $\mathbf{x}^{l}_{v}=\psi\bigg{(}\mathbf{x}^{l-1}_{v},\phi(\mathbf{x}^{l-1}_{u}:\mathcal{A}(v,u)=1)\bigg{)}$
Where $\phi$ is an aggregation function, and $\psi$ is a combination function, typically permutation-invariant operations like Sum, Mean, or Max. $\mathbf{x}^{l}_{v}$ and $\mathbf{x}^{l-1}_{v}\in\mathbb{R}^{d_{hidden}}$ represent the output of node $v$ from layers $l$ and $l-1$, respectively. $\mathbf{x}^{l-1}_{u}\in\mathbb{R}^{d_{hidden}}$ represents the output of $v$’s neighbor at the previous layer. The final layer representation $\mathbf{x}^{k}_{v}$ is then passed through a Multilayer Perceptron (MLP) to reduces the dimensionality of the representation to $d_{out}$. By altering GNNs with the ones considering edge information, this method can be extended to edge-labeled graphs without losing generalization.
4.2. Node Matching Cost Estimation

### v2

# Flexible Graph Similarity Computation With A Proactive Optimization Strategy

@@ Abstract
Graph Edit Distance (GED) offers a principled and flexible measure of graph similarity, as it quantifies the minimum cost needed to transform one graph into another with customizable edit operation costs. Despite recent learning-based efforts to approximate GED via vector space representations, existing methods struggle with adapting to varying operation costs. Furthermore, they suffer from inefficient, reactive mapping refinements due to reliance on isolated node-level distance as guidance. To address these issues, we propose Gen, a novel learning-based approach for flexible GED approximation. Gen addresses the varying costs adaptation by integrating operation costs prior to match establishment, enabling mappings to dynamically adapt to cost variations. Furthermore, Gen introduces a proactive guidance optimization strategy that captures graph-level dependencies between matches, allowing informed matching decisions in a single step without costly iterative refinements. Extensive evaluations on real-world and synthetic datasets demonstrate that Gen achieves up to 37.8% reduction in GED approximation error and 72.7% reduction in inference time compared with state-of-the-art methods, while consistently maintaining robustness under diverse cost settings and graph sizes.
1 Introduction
@@
Measuring graph similarity is a fundamental problem in graph mining, with wide-ranging applications in retrieval, classification, and recommendation tasks [8, 11, 18, 27, 32, 33, 34]. Among various similarity measures, Graph Edit Distance (GED) is particularly appealing due to its flexibility: it defines the similarity between two graphs as the minimum cost of transforming one into another through node/edge insertion, deletion, or substitution, under customizable cost settings. This flexibility allows GED to be tailored to specific applications. For example, in graph databases for recommender systems, higher costs can be assigned to nodes/edges representing key user preferences. This prioritizes the retrieval of graphs that preserve these preferences, enhancing personalized recommendations.
Traditionally, exact GED computation relies on A*-like solvers [2, 8, 30], which sequentially construct node/edge mappings. In this procedure, operation costs critically influence matching decisions, resulting in varying optimal mappings under different cost settings, as illustrated in Figure 1. Moreover, the sequential manner maintains structural consistency with prior decisions, and inherently models dependencies among node matches induced by intra-graph connections and inter-graph discrepancies. Such a procedure effectively embodies two fundamental principles of GED computation: the influence of operation costs and the dependencies between node matches. However, exhaustive search incurs prohibitive computational costs. Even for graphs with as few as 16 nodes, exact GED computation remains infeasible within practical time limits [6]. This bottleneck severely restricts the applicability of traditional methods in time-sensitive graph retrieval scenarios.
In response, recent advances leverage Graph Neural Networks (GNNs) [7, 15, 26, 28] to approximate GED by embedding graphs into vector spaces [3, 4, 5, 9, 13, 14, 16, 17, 20, 21, 22, 31, 35]. Existing learning-based methods generally fall into two categories: Graph comparison approaches compress node representations into holistic graph embeddings. In contrast, node comparison approaches estimate edit costs between node representations and iteratively refine mappings by down-weighting high-cost matches. Although graph comparison methods offer computational efficiency, they lose critical node-level details essential for modeling edit operations. In contrast, node comparison methods retain finer structural information, making them more suited to capture the flexibility of GED.
Figure 1: The black dashed lines are matches, and the red markings are edits. $f_{1}$ and $f_{2}$ are mappings. $f_{1}$ incurs 2 node substitutions (ns), 1 node insertion, 1 edge deletion, and 2 edge insertions (ei), denoted as $[2,1,0,1,2]$, while $f_{2}$ incurs $[4,1,0,0,1]$. When ns costs 2, other operations cost 1, $f_{1}$ is optimal with a cost of 8, while $f_{2}$ costs 10. However, if ei costs 2, others cost 1, $f_{1}$ costs 8 while $f_{2}$ costs 7, making $f_{2}$ the optimal choice.
Nevertheless, existing node comparison approaches exhibit two core limitations that directly violate the above principles of GED computation. First, they neglect the impact of operation costs during mapping determination. While GraphEdX [13] incorporates these costs when computing the final GED, it ignores them during mapping establishment, resulting in static mappings that do not adapt to cost variations. Second, these methods rely on node-level similarity scores to guide matching decisions, ignoring the graph-level dependencies between node matches, which reflect how matches jointly affect global structural alignment. Without modeling them, mappings are optimized in isolation, often resulting in myopic matches and necessitating costly post hoc refinement procedures, which substantially increase computational overhead.
In this paper, we propose Graph Edit Network (Gen), a novel framework that addresses these limitations by revisiting the fundamental principles of GED computation. Gen introduces two key innovations: (1) it incorporates operation costs prior to mapping establishment, ensuring that mapping decisions dynamically adapt to cost variations, and (2) it explicitly models graph-level dependencies between matches through a proactive guidance optimization strategy. By globally estimating and propagating node alignment difficulties across graphs, Gen captures the joint impact of individual node matches, facilitating informed one-shot matching and eliminating the need for costly refinements.
The contributions of this work are threefold: (i) We introduce a principled yet simple approach that integrates operation costs prior to mapping determination, addressing a fundamental gap in learning-based graph similarity models and improving adaptability to varying operation cost settings. (ii) We propose a proactive guidance optimization strategy, utilizing global alignment difficulty estimation and difficulty propagation to capture graph-level matching dependencies, enabling accurate one-shot matching without iterative refinements. (iii) Extensive empirical evaluations demonstrate the effectiveness, robustness of Gen, achieving substantial improvements in both accuracy and efficiency over existing methods.
2 Related Work
Graph similarity computation (GSC) aims to approximate measures such as graph edit distance (GED) [4, 16] using neural network-based approaches. This work focuses on end-to-end GED approximation methods, which can be further divided into graph-level and node-level comparison approaches. For more details, we refer the reader to recent surveys on this topic [18].
Graph-level comparison methods embed entire graphs into latent spaces and compute similarity based on graph representations [16, 21, 35, 22, 3]. GMN-emb [16] employs a gated mechanism to highlight important nodes during graph embedding. EGSC [21] introduces a teacher-student framework to balance prediction accuracy and storage efficiency. Eric [35] enforces node-graph alignment during training, and detaches it to accelerate inference. Greed [22] models the triangle inequality among graph embeddings by leveraging bounds on GED. To improve interpretability, TaGSim [3] predicts graph edit vectors (GEVs) representing edit operation counts. However, graph-level comparison methods lack sensitivity to mapping variations, which may result in limited accuracy under varying cost settings.
@@ 2 Related Work
Node-level comparison methods compares node embeddings to capture fine-grained node similarities [16, 4, 5, 9, 20, 13]. SimGNN [4] augments graph representations with non-differentiable histogram features derived from node similarities, which cannot be directly learned during the training process. GMN-match [16] iteratively fuses cross-graph node representations using distance-based local attention, which may lead to suboptimal, myopic updates. GraphSim [5] reorders multi-scale similarity matrices via BFS traversal, which may oversimplify complex structures and reduce robustness. Recent approaches introduce combinatorial optimization to refine mappings. GOTSim [9] augments the similarity matrix to account for node deletion and insertion operations, then uses a linear assignment solver to align nodes. GEDGNN [20] and GraphEdX [13] leverage Gumbel-Sinkhorn networks for differentiable mapping, with GraphEdX further designing operation-specific distance measures. Nevertheless, these methods primarily operate in matrix form and often ignore graph-level matching dependencies, which may lead to suboptimal mappings that fail to reflect the true graph structural differences. Moreover, all existing methods compute static mappings independent of operation costs, limiting their adaptability to varying cost settings.
3 Background
@@
Graph Edit Distance (GED). Given a source graph $S$ and a target graph $T$, the graph with fewer nodes was padded using dummy nodes to ensure $|\mathcal{V}_{S}|=|\mathcal{V}_{T}|$. GED is defined as the minimum total edit cost over all bijective mappings $f\in\mathcal{B}:\mathcal{V}_{S}\mapsto\mathcal{V}_{T}$, where the bipartite graph $\mathcal{B}=(\mathcal{V}_{S}\cup\mathcal{V}_{T},\mathcal{E}_{\mathcal{B}})$ represents the solution space of $f$, and each edge $m\in\mathcal{E}_{\mathcal{B}}$ denotes a potential node correspondence between $S$ and $T$. The GED from $S$ to $T$ is computed as: